The angles of a polygon add up to 2700 degrees. How many sides does the polygon have?
Answer by venugopalramana(3286)

The angles of a polygon add up to 2700 degrees. How many sides does the polygon have?
SUM OF ANGLES =(2N-4)*90 DEGRES=2700
2N-4=2700/90=30
2N=30+4=34
N=34/2=17